It's (3b over 4) minus (1 over 5) = (b over 2) i wanna solve and make b the subject. I would really appreciate step by step working out thanks :)

Answer by checkley79(3341)

3B/4-1/5=B/2
3B/4-B/2=1/5
(3B-2*B)/4=1/5 CROSS MULTIPLY.
(3B-2B)/4=1/5
B/4=1/5 CROSS MULTIPLY.
5B=4
B=4/5 ANS.
PROOF:
3(4/5)/4-1/5=(4/5)/2
(12/5)/4-1/5=4/5*1/2
12/5*1/4-1/5=4/10
12/20-1/5=2/5
3/5-1/5=2/5
2/5=2/5